The angle between two straight line 2x=3y=-z and 6x=-y=-4z

Grade:12

1 Answers

Arun
6 years ago
Dear Anand
 
 
Angle between the lines 2x = 3y = - z & 6x = -y = -4z is 90°
Since a1a2 + b1b2 + c1c2 = 0.
where a1=1/2,a2=1/6,b1=1/3,b2=-1,c1=-1,c2=-1/4
 
Thanks and Regards
Arun (askIITians forum expert)